Please help with this Kinematics question! Position and time graphs. Velocity vs. time graph.
Consider this position vs. does the corresponding velocity vs. time graph look like? 1. time graph. What (a) Sketch your prediction below * Now reproduce the position graph by walking, and see if the velocity graph comes out the way you expect. 10 t (seconds) Velocity Graph Prediction Velocity Graph Result (m/s) o (m/s) 0 10 6 10 t (seconds) t (seconds) 2. Heres a velocity vs. time graph. What does the corresponding position graph look like? (a) Sketch your prediction below *Now reproduce the velocity graph by walking, and (m/s) 0 :.t(seconds) see if the position graph comes out the way you expect Postition Graph Prediction Position Graph Result t (seconds) t (seconds)
Answer #1

(1)

Slope of Position vs time graph gives Velocity vs time graph.

From t=0 to t=5 s, Slope is decreasing with time but slope is positive

i.e velocity is positive but it is decreasing with time

After t=5 s, slope becomes negative

From t=5 s to t=10 s, Slope is negative at all points and velocity is decreasing

Following is the graph

个. ? 那-S

(2)

As we know,

Area under the curve velocity vs time graph gives displacement.

and .

slope of Position vs time graph gives Velocity vs time graph

Area under curve from t=0 to t=3 s, is A= 3*1/2 = 1.5 m

Since velocity is positive constant, thus shape of curve in position vs time graph must be straight line with positive slope.

Area under curve from t=3 s to t=5 s, is A= 2*1/2 = 1 m

Since velocity is straight line with negative slope, thus shape of curve in position vs time graph must be parabolic curve opening downwards.

from t=5 s to t=7 s, velocity is zero, i.e position remains same.

From t=7 s to 10 s, Area under curve, A = 3*-1/2 = -1.5 m

Since velocity is negative constant, thus shape of curve in position vs time graph must be straight line with negative slope.

Following is the graph